A student heated 21.5 g pure hydrated calcium sulfate ,CaSO4 x H2O . 17.0 g anhydrous calcium sulfate was formed .Determine the molecular formula for this hydrated salt
1
Expert's answer
2019-10-09T07:36:35-0400

In 17 g CaSO4, (21.5-17)=4.5 g H2O was present.

Molecular weight of CaSO4=136g/mole

Molecular weight of H2O = 18 g/mole

17 g CaSO4 = 17/136 moles

=0.125 moles

4.5 g H20 = 4.5/18 moles

= 0.25 moles

In 0.125 moles CaSO4 there are 0.25 moles H2O

In 1 mole CaSO4 there are 0.25/0.125 mole of ahH2O

=2 moles of H2O

Therefore the correct formula is = CaSO4, 2H2O